SpaceRef

SpaceRef


Aqua Weekly Report 7-14 June 2002

Status Report From: Goddard Space Flight Center
Posted: Friday, June 14, 2002

Checkout of the Aqua spacecraft continues to go extremely well. We completed our third, fourth, and fifth Ascent Maneuvers nominally on Saturday (6/8), Monday (6/10), and Thursday (6/13), respectively. The final Ascent Maneuver is now scheduled for next Monday (6/17).

The spacecraft experienced its first lunar eclipse of the sun on Monday (6/10). The eclipse duration was 13 minutes, reaching a maximum shadowing of 25%. EPS supermode was set to "survival" prior to the eclipse to achieve maximum charging and the peak DOD fault response was disabled. The battery was fully charged going into the eclipse and remained so throughout the event. A reduction in solar array panel temperature was noted during the eclipse. A second lunar eclipse, of lesser duration and shadowing, occurred on the next orbit.

Instrument checkout activities also continued nominally throughout the week. The instrument status is:

AIRS: Currently in Checkout mode. AIRS turned off their decontaminate heater and initialized their cryocoolers on Monday (6/10). Wednesday (6/12), AIRS began operating with their cryocoolers, transitioned to Ready mode, entered their Calibration state, and ran their first Run Guard Test. They also have begun executing activities out of the MCL, including subsequent (daily) Run Guard Tests. Further checkout activities are planned for next week.

AMSR-E: Currently in Normal mode, rotating at 40 rpm. AMSR-E has completed their activation and checkout activities. No further instances of the SPS anomaly occurred last week and the investigation into the cause of the anomaly continues.

AMSU-A: Currently in Full Scan mode. Other than parking the instruments for the Ascent Maneuvers, AMSU had no scheduled activities this past week. They plan to perform their Space View Calibrations next week after we complete the final Ascent Maneuver.

CERES: Currently in Aft-Crosstrack, Fore-Biaxial mode with their covers closed. Other than stopping scanning for the Ascent Maneuvers, CERES had no scheduled activities this past week. They plan to open their covers next week after we complete the final Ascent Maneuver.

HSB: Currently in Scan Normal mode. Other than parking the instrument for the Ascent Maneuvers, HSB had no scheduled activities this past week. They plan to enter Investigation mode next week after we complete the final Ascent Maneuver.

MODIS: Currently in Science mode, with the Nadir Aperture Door (NAD) and Solar Diffuser Door (SDD) unlatched and closed, and the Space View Door (SVD) unlatched and at the OPEN position. MODIS transitioned to science mode and performed their first Photovoltaic Electronic Calibration (PV ECal) last Friday (6/7). On Tuesday (6/11), MODIS performed several activities, including a Sector Rotation, testing the SDD operation, executing a functional test of the Solar Diffuser Stability Monitor (SDSM), dumping the Door Table Memory and confirming proper operation, and performing another PV ECal (via the MCL). Today (6/14), the first set of six MODIS Yaw maneuvers will be performed, with the SDD screened and the NAD closed. They plan additional Yaw maneuvers and further checkout activities next week.

With the MODIS transition to science mode on Friday (6/7), the final source of high rate science data began flowing into the Solid State Recorder (SSR). The SSR is now recording data at ~10.36 Mbps (was ~1.73 Mbps). To date, 100% of the recorded data has been successfully dumped to the EPGN/EDOS.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.